What is the velocity of a shot put with a mass of 7.3 kg and momentum of 22 kg -m/s moving in the forward direction?


1
Expert's answer
2021-12-01T10:28:43-0500
"p=mv,""v=\dfrac{p}{m}=\dfrac{22\ \dfrac{kg\times m}{s}}{7.3\ kg}=3\ \dfrac{m}{s}."